A railroad car is moving along a straight frictionless track. In each of the following cases, the car initally has a total mass (car and contents) of 200kg and is travelling with a velocity 4m/s. Find the final velocity of the car in each of the 2 cases:                                                                      (a) A 20Kg mass is thrown sideways out of the car with a velocity of 2m/s relative to the car?
(b) A 20kg mass is thrown into the car with a velocity of 6m/s relative to the ground and opposite in direction to the velocity of the car ?                                        
 
 

Grade:11

1 Answers

Manish
5 years ago
(a)800=200v1
v1=800/200=4m/s
(b) P=mv =200*4=800
P(final)=m1v1+m2v2
800=200v1+20*(-4)
v1=880/200
v1=4.4m/s